OpenStudy (anonymous):

Find the inverse function of y=(4x-1)/(2x+3)

hartnn (hartnn):

do you know the steps to find inverse function ? have you tried this one ?

OpenStudy (anonymous):

I know the steps, but i'm confused as to whether to simply multiply 2x+3 across or not

hartnn (hartnn):

yes, you should do that .....and tell me what u get ?

OpenStudy (anonymous):

[(2x+4)*(Y)]/4=x

hartnn (hartnn):

how did 2x+3 became 2x+4 ?

OpenStudy (anonymous):

I brought the 1 over from the 4x-1 equation

OpenStudy (anonymous):

added 1 to both sides.

hartnn (hartnn):

the 1st step is to multiply 2x+3 (2x+3) y = 4x-1 got this ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

good, now distribute the y (2x+3) y= ... ?

OpenStudy (anonymous):

2xy+3y

OpenStudy (anonymous):

then I move the 1 over, and divide by 4?

hartnn (hartnn):

2xy+3y = 4x-1 the goal is to isolate 'x' , so bring all 'x' terms together for that you'll either need to bring 2xy on right or 4x on left do whatever you are comfortable with...

hartnn (hartnn):

remember first bring all 'x' terms together

OpenStudy (anonymous):

4x-2xy-1=3y

OpenStudy (anonymous):

then, (4x-2xy-1)/3=y?

hartnn (hartnn):

correct, also get rid of the 1 from left first so, after 4x-2xy-1 = 3y you would add 1 to both sides

OpenStudy (anonymous):

What does that serve to accomplish? I thought the inverse function was to be expressed in terms of y?

hartnn (hartnn):

in terms of 'y' , after you interchange 'x' and 'y'

OpenStudy (anonymous):

ah, I had overlooked that. Thank you very much

hartnn (hartnn):

the main point is to isolate other variable

hartnn (hartnn):

lets continue ? 4x-2xy-1 = 3y then ?

OpenStudy (anonymous):

4x-2xy=3y+1

OpenStudy (anonymous):

Can we simplify left to be 2x(2-y)=3y+1?

hartnn (hartnn):

correct!

hartnn (hartnn):

now you can just divide both sides by 2 (2-y) and you would have isolated x :)

OpenStudy (anonymous):

Then divide right side by 2(2-y), to isolate x variable?

OpenStudy (anonymous):

Awesome, thank you very much for the professional walkthrough!

hartnn (hartnn):

you got it :)

hartnn (hartnn):

welcome ^_^